(CH16) Part A) The pOH of an aqueous solution of 0.377 M hydroxylamine (a weak base with the form...

(CH16)

Part A) The pOH of an aqueous solution of 0.377 M hydroxylamine (a weak base with the formula NH2OH) is

Part B) The hydronium ion concentration of an aqueous solution of 0.504 M codeine (a weak base with the formula C18H21O3N) is ...

[H3O+] =  M.
Answer #1

part A:

pKb = 7.97

pOH = 1/2 * [PKb - log C]

or

pOH = 0.5 * [7.97 - log (0.377)]

or

pOH = 4.20

part B:

pKb = 5.79.

pOH = 1/2 * [PKb - log C]

or

pOH = 0.5 * [5.79 - log (0.504)]

or

pOH = 3.04

pH = 14 - 3.04 = 10.96

-log [H3O+] = 10.96

or

[H3O+] = 1.11 * 10^-11 M